### v1

@@ Abstract
With the rapid advancement of high-power computing and electronic technologies, understanding thermal conductivity in materials has become increasingly important. This study presents a novel method called the Space-Projected Thermal Conductivity (SPTC) that can predict thermal activity at the atomic scale, leveraging the Green-Kubo thermal transport equations. We demonstrate the effectiveness of this approach on crystalline and amorphous carbon-based structures, paracrystalline and amorphous silicon, and grain boundaries in silicon-germanium alloys. The results highlight the potential of our method to provide new insights into the thermal behavior of materials, offering a promising avenue for materials design and performance optimization.
@@ 1 Introduction
In this study, we present the Space-Projected Thermal Conductivity (SPTC) method, which offers an intuitive framework for identifying thermally active sites within materials in real space. We adopt the Allen and Feldman approach [12] to derive an expression for thermal conductivity using the Green-Kubo formalism for thermal transport [13]. The SPTC method is applied to a range of crystalline, defective, and amorphous materials, including carbon and silicon allotropes. Additionally, we investigate the thermal properties of interfaces and grain boundaries in paracrystalline silicon and silicon-germanium alloy systems. It is important to note that this method accounts solely for phonon contributions to heat transport, excluding electronic effects.
@@ 2 Theory
$\mathbf{S}=-\frac{1}{2V}\sum_{i,j}\sum_{\alpha,\beta}({\bf R}_{i}-{\bf R}_{j})\frac{p_{i\alpha}}{m_{i}}\frac{\partial^{2}E}{{\partial u_{i}^{\alpha}\partial u_{j}^{\beta}}}u_{j}^{\beta}.$ (2)
where, ${\bf R}_{i}$ in the position of ith atom in the unit cell. Invoking the usual harmonic approximation (HA), the classical Hamiltonian is:
$\mathcal{H}=\sum_{i}\frac{p_{i}^{2}}{2m_{i}}+\frac{1}{2}\sum_{i,j}\frac{\partial^{2}E}{\partial u_{i}\partial u_{j}}u_{i}u_{j},$ (3)
where $p_{i}$ and $u_{i}$ are the momenta and small displacements. The potential energy, $E$, is given as:
$E=\frac{1}{2}\sum_{\gamma,\gamma^{\prime}}\sum_{i,j}\sum_{\alpha,\beta}\phi_{i,j}^{\alpha\beta}(\gamma,\gamma^{\prime})u_{\gamma i}^{\alpha}u_{\gamma^{\prime}j}^{\beta}$ (4a)
$\phi_{i,j}^{\alpha\beta}(\gamma,\gamma^{\prime})=\frac{\partial^{2}E}{\partial u_{\gamma i}^{\alpha}\partial u_{\gamma^{\prime}j}^{\beta}}.$ (4b)
@@
$\omega^{2}u_{i}^{\alpha,{\bf k}}$ $=\sum_{\beta j}D^{\alpha\beta}_{ij}\big{(}{\bf k}\big{)}u_{j}^{\beta,{\bf k}},$ (5a)
@@
$D^{\alpha\beta}_{ij}({\bf k})$ is the Dynamical Matrix (DM), the Fourier transform of the force constant matrix. ${\bf R}_{j,\gamma}$ is the position of $j^{th}$ atom in the $\gamma^{th}$ cell. k is a wave vector that may be chosen at the center or corner of the Brillouin zone. For $N$ cells in a system with $l$ atoms per cell, labeled by $x$ = 1, 2 . . . , $l$; the S operator from Equation 2 is transformed to:
@@
Here, $m,n$ are the indices of the phonon modes; $\hat{a}_{m}^{\dagger}$ and $\hat{a}_{n}$ are the creation and annihilation operators of a phonon mode respectively; ${\bf v}^{\bf k}_{mn}$ is the group velocity vector of mode $m$ when $m=n$; $\omega_{m}^{\bf k}$ is the vibrational frequency of the mode $m$. Upon applying the hamonic appoximation results on Equation 1, a general expression for the real part of TC is obtained [12]:
@@
where, $f_{m}$ is the equilibrium occupation of the $m^{th}$ mode (i.e., the Bose-Einstein phonon distribution); $(S^{\alpha})_{mn}$ is the $\alpha^{th}$ component of $S_{mn}$. When $\omega\rightarrow 0$, due to the presence of the $\delta$ function, the term $\frac{\big{<}f_{m}\big{>}-\big{<}f_{n}\big{>}}{\hbar(\omega_{m}-\omega_{n})}$ in Equation 7 simplifies to $-{\partial\big{<}f_{m}\big{>}}/{\partial\omega_{m}}$. With these considerations, and using Equations 6, Allen and Feldman’s formulation for TC is expressed as:
@@
where, ${R}_{\gamma}^{\eta}$ is the $\eta$ component of position vector for cell $\gamma$; ${R}_{xx^{\prime}}^{\eta}$ is the $\eta$ component of displacement between atoms $x$ and $x^{\prime}$; $e^{\alpha,m}_{x,{\bf k}}$ is the polarization of the vibrational mode $m$ at atom $x$ along $\alpha$ direction. The thermal conductivity in Equation 2, is taken as an average of the Cartesian components of the conductivity tensor. In the case of $\textbf{{k}}={\textbf{{0}}}$, Equation 2 is reorganized as:
$\kappa=$ $\frac{1}{V}\sum_{\eta}\sum_{\gamma,x,x^{\prime}}\frac{1}{\sqrt{m_{x^{\prime}}m_{x}}}\left({R}^{\eta}_{\gamma}+{R}^{\eta}_{xx^{\prime}}\right)\sum_{m,n\neq m}\delta(\omega^{\bf{0}}_{m}-\omega^{\bf{0}}_{n})$
$\frac{\pi V^{2}}{3\hbar^{2}\omega_{m}^{2}}\left(\frac{\hbar}{2V}\right)^{2}(\omega^{\bf{0}}_{m}+\omega^{\bf{0}}_{n})^{2}\frac{1}{4\omega^{\bf{0}}_{m}\omega^{\bf{0}}_{n}}\frac{\hbar\omega_{m}^{2}}{T}$
$\left(\frac{\hbar}{k_{B}T}\frac{e^{\frac{\hbar\omega_{m}}{k_{B}T}}}{\left(e^{\frac{\hbar\omega_{m}}{k_{B}T}}-1\right)^{2}}\right)\sum_{\alpha,\beta}\phi^{\alpha\beta}_{x,x^{\prime}}({0},\gamma)e^{\alpha,m}_{x,\bf{0}}e^{\beta,n}_{x^{\prime},\bf{0}}$
$\sum_{\gamma^{\prime},a,b}\sum_{\alpha^{\prime},\beta^{\prime}}\frac{1}{\sqrt{m_{a}m_{b}}}\phi^{\alpha^{\prime}\beta^{\prime}}_{a,b}({0},\gamma^{\prime})e^{\alpha^{\prime},m\dagger}_{a,\bf{0}}e^{\beta^{\prime},n\dagger}_{b,\bf{0}}\left(R^{\eta}_{\gamma^{\prime}}+R^{\eta}_{ab}\right)$ (9)
2.1 The Space-Projected Thermal Conductivity Formalism
@@
Rearranging summation order in Equation 2, we can obtain an expression for $\kappa$ as a (double) sum over terms depending on atomic positions $x$ and $x^{\prime}$:
@@
For $\omega=0$, $\Xi$ is a Hermitian matrix, i.e., $\Xi(x,x^{\prime})=\Xi(x^{\prime},x)$. Equation 10 suggests a way to decompose the total TC into contributions depending on atomic position $x$ by conduction a summation of $\Xi(x,x^{\prime})$ over all positions $x^{\prime}$:
@@
Here, $\zeta(x)$ can be named the space-projected thermal conductivity (SPTC), as it represents the contribution of atom at site $x$ to the total TC of the system. Summing $\zeta(x)$ over atom index $x$ recovers the total TC $\kappa$:
@@ 3 Illustration of the Method
As a proof of concept, we conducted SPTC analysis on a range of different material systems, primarily carbon, silicon, and silicon-germanium structures. These systems included crystalline graphene, with both single and double vacancy defects, as well as amorphous graphene, characterized by ring disorder. Additionally, we applied SPTC analysis to diamond-crystalline silicon, paracrystalline silicon, and silicon-germanium alloy systems, with a focus on grain boundaries and mass defects.
Molecular dynamics (MD) simulations were conducted using Large-scale Atomic/Molecular Massively Parallel Simulator (LAMMPS) [14]. Within LAMMPS, the Machine Learning Gaussian approximation potential (ML-GAP) for carbon [15] and silicon [16] were employed for the carbon and pure silicon simulations. The Tersoff potential was used for the silicon-germanium structure [17].
@@ 3.1 Carbon Structures
Figure 1: Representations of the carbon-based models: (a) crystalline graphene, (b) 585 Stone-Wales defect, (c) single vacancy defect, and (d) the undulating sp2 structure of amorphous graphene.
@@ 3.2 Silicon Single-Vacancy Defect
A diamond-structure Si unit cell was sourced from the Materials Project [26], and a 1000-atom, $5\times 5\times 5$ supercell model was constructed from it and used as the base for creating different Si structures. The original cubic supercell, with length of 27.21 Å  and density ($\rho$) of 2.33 g/cm-3, was relaxed, exhibiting no significant change in its conformation. A single-vacancy defect, shown in Figure 2(a), was created by removing an atom from the crystalline Si lattice, followed by another energy/structural relaxation. For the analysis of the temperature and frequency dependence of TC in amorphous silicon (aSi), a tetrahedral Si model with 512 atoms ($\rho=2.34$ g/cm-3), obtained from Reference [27], was utilized.
Figure 2: Representations of the silicon models (yellow spheres): (a) single-vacancy defect in diamond Si, (b) $\Sigma 3(112)/[1\bar{1}0]$ twin grain boundary, with five-fold coordinated Si atoms at the boundary shown as dark green spheres, and (c) silicon-germanium structure, where germanium atoms (cyan spheres) are randomly substituted for Si atoms in the diamond Si lattice.
@@ 4.1 Carbon Structures
The SPTC for three-fold coordinated hexagonal rings in crystalline graphene are uniformly distributed. However, introducing a single-atom vacancy in an otherwise crystalline graphene structure significantly reduces the TC of nearby atoms. The values for the two-fold coordinated carbon atoms near the vacancy is approximately 27% of that for the three-fold coordinated atoms in the bulk region.
As shown in Figure 3(a), the atoms adjacent to the vacancy (highlighted in blue) exhibit the lowest SPTC. An incremental, radially isotropic increase in normalized SPTC is observed moving outward from the defect center. To quantify anisotropy, the SPTC along two angular directions — 0 and 30, marked by the red and green lines in Figure 3(a) — is shown in Figure 3(b). The variation in SPTC with distance follows a similar trend along both directions.
Figure 3: (a) Plot of the normalized SPTC results for a single vacancy in crystalline graphene. The incremental changes in SPTC values, originating from the defect center and radiating outward are plotted in (b). Likewise, the SPTC and their incremental changes from the defect center in the SW-585 crystalline graphene configuration are depicted in (c) and (d), respectively.
@@
In the defective SW-585 configuration shown in Figure 3(c), the atoms in the five- and eight-membered rings are the lowest thermal conduction active sites, with the normalized SPTC reduced by approximately 60% compared to the bulk maximum. The SPTC gradually increases in all directions moving away from the defective region. The changes along two angular directions, 90 and 30, represented by red and green lines in Figure 3(c), are illustrated in Figure 3(d). The plot shows isotropic increment in SPTC, returning to the maximum (bulk) values at around 10 Å  away from the defect center.
@@
In the aGr2000 model shown in Figure 4[i], atoms with only two neighbors exhibit SPTC values up to 74% of those of three-coordinated atoms. However, the SPTC for the three-coordinated carbon atoms vary depending on their local ring connectivity: atoms in non-hexagonal rings are less thermally active than those in hexagonal rings, similar to what is observed in crystalline graphene. To confirm this, a 60-atom amorphous graphene model (aGr60) was created. Its SPTC analysis, shown in Figure 4[ii], also indicates higher values for hexagonal rings compared to non-hexagonal ones. The ring distribution for both amorphous graphene models is illustrated in Figure 4[iii].
We note a similarity between thermally and electronically conductive regions. References [19] and [20], which report the electronic transport properties of amorphous graphite, report that in the amorphous graphene layers, the hexagonal carbon rings favor electron flow, while connections between pentagonal and heptagonal rings impede it. Similar observation was made for ring disorder in amorphous carbon nanotubes [29].
Figure 4: Normalized SPTC results for amorphous graphene with [i] 2000 atoms (aGr2000) and [ii] 60 atoms (aGr60). The ring distribution for the aGr2000 and aGr60 models is shown in [iii].
@@ 4.2 Silicon Structures
To test our implementation of Allen and Feldman formulation for the thermal conductivity [12], we compared the temperature and frequency dependence of TC in amorphous silicon produced by our implementation with the results of Feldman et al. [30], as shown in Figure 5(a) and (b). The TC temperature dependence obtained here is consistent with that in Reference [30], while the peak in the frequency dependence curve (the gray dashed line at 0.24 meV) matches the position they reported at 175 K. Note that the consistency is strong despite the use of a diiferent interatomic potential and model in our work.
Figure 5: (a) Temperature dependence of TC for amorphous silicon, with results from this study compared to those of Feldman and coworkers [30]. (b) Frequency dependence of TC at various temperatures, with the dashed gray line marking the peak at 175 K. This peak aligns with the corresponding plot at the same temperature in Reference [30].
@@
In defect-free diamond-silicon structure, all atoms contribute equally to the total TC of the system. On the other hand, local TC on atomic sites in the vicinity of the single-vacancy defect is reduced. Such under-coordinated silicon atoms with dangling bonds exhibit $\approx$70reduction in SPTC, as compared to the the bulk-crystal value. The SPTC evolution with distance from the center of the defect in the (111) plane — along green and red lines shown in Figure 6(a) — is presented in Figure 6(b).
Figure 6: (a) Normalized SPTC data for a single-vacancy defect in crystalline Si, projected along the (111) plane intersecting the defect site. (b) The variation in SPTC with distance from the defect center along the red and green lines shown in panel (a).
@@
In the SPTC analysis of the $\Sigma 3(112)/[1\bar{1}0]$ silicon grain boundary in Figure 7(a), the five-coordinated atoms (floating bonds) at the grain boundary exhibit lower SPTC compared to the bulk silicon and also influences the thermal activity of nearby atomic sites. Futhermore, a strong correlation between bonding environment and SPTC is evident: in Figure 7(b), the five-coordinated Si atoms labeled “i” show an SPTC reduction of around 79% compared to the bulk crystalline region labeled “ii.” Figure 7(c) further indicates that bond length distortions at the grain boundary contribute to the reduced SPTC, as atomic pairs with bond lengths exceeding 2.36 Å (the bond length in diamond silicon) exhibit significantly lower SPTC values.
Figure 7: (a) SPTC for the Si $\Sigma 3(112)/[1\bar{1}0]$ twin grain boundary (GB), highlighting the correlation between SPTC values and bond lengths in different Si configurations: (i) five-fold coordination at the GB and (ii) four-fold coordination at the bulk region. The variation in SPTC for atoms along the line labelled “i” and “ii” is shown in (b), with their bond length distributions depicted in (c).

### v2

@@ Abstract
With the rapid advance of high-performance computing and electronic technologies, understanding thermal conductivity in materials has become increasingly important. This study presents a novel method: the Site-projected Thermal Conductivity (SPTC) that quantitatively estimates the local (atomic) contribution to heat transport, leveraging the Green-Kubo thermal transport equations. We demonstrate the effectiveness of this approach on disordered and amorphous graphene, amorphous silicon, and grain boundaries in silicon-germanium alloys. Amorphous graphene reveals a percolation behavior for thermal transport. The results highlight the potential of our method to provide new insights into the thermal behavior of materials, offering a promising avenue for materials design and performance optimization.
keywords:
thermal conductivity, defects, molecular simulation, green-kubo equation, SPTC
A. Gautam Y. G. Lee C. Ugwumadu K. Nepal S. Nakhmanson D. A. Drabold*
{affiliations}
A. Gautam, Y. G. Lee, K. Nepal
Department of Physics and Astronomy, Ohio University, Athens, 45701, OH, USA
Dr. C. Ugwumadu
Physics of Condensed Matter and Complex Systems (T-4) Group, Los Alamos National Laboratory, Los Alamos, 87545, NM, USA
Prof. S. Nakhmanson
Department of Physics, Department of Materials Science & Engineering, Institute of Materials Science, University of Connecticut, Storrs, 06269, CT, USA
Prof. D. A. Drabold
Department of Physics and Astronomy, Ohio University, Athens, 45701, OH, USA
drabold@ohio.edu
@@ 1 Introduction
In this study, we present the Site-Projected Thermal Conductivity (SPTC) method, which offers an intuitive framework for identifying thermally active sites within materials in real space. We adopt the Allen and Feldman approach [12] to derive an expression for thermal conductivity using the Green-Kubo formalism for thermal transport [13]. The method is formulated for disordered systems including topological and chemical disorders, extended, and point defects. The SPTC method is applied to a range of crystalline, defective, and amorphous materials, including carbon and silicon allotropes. Additionally, we investigate the thermal properties of interfaces and grain boundaries in paracrystalline silicon and silicon-germanium alloy systems. It is important to note that this method accounts solely for phonon contributions to heat transport, excluding electronic effects.
@@ 2 Theory
$\mathbf{S}=-\frac{1}{2V}\sum_{i,j}\sum_{\alpha,\beta}({\bf R}_{i}-{\bf R}_{j})\frac{p_{i\alpha}}{m_{i}}\frac{\partial^{2}U}{{\partial q_{i}^{\alpha}\partial q_{j}^{\beta}}}q_{j}^{\beta}$ (2)
where, ${\bf R}_{i}$ in the position of $i^{th}$ atom in the unit cell, and $q_{i}^{\alpha}$ is the $\alpha$-component of the small displacement vector for $i^{th}$ atom. Invoking the usual harmonic approximation (HA), the potential energy, $U$, is given as:
$U=\frac{1}{2}\sum_{\gamma,\gamma^{\prime}}\sum_{i,j}\sum_{\alpha,\beta}\phi_{i,j}^{\alpha\beta}(\gamma,\gamma^{\prime})q_{\gamma i}^{\alpha}q_{\gamma^{\prime}j}^{\beta}$ (3a)
$\phi_{i,j}^{\alpha\beta}(\gamma,\gamma^{\prime})=\frac{\partial^{2}U}{\partial q_{\gamma i}^{\alpha}\partial q_{\gamma^{\prime}j}^{\beta}}.$ (3b)
@@
$\omega^{2}q_{i}^{\alpha,{\bf k}}$ $=\sum_{\beta j}D^{\alpha\beta}_{ij}\big{(}{\bf k}\big{)}q_{j}^{\beta,{\bf k}},$ (4a)
@@
$D^{\alpha\beta}_{ij}({\bf k})$ is the Dynamical Matrix (DM), the Fourier transform of the force constant matrix. ${\bf R}_{j,\gamma}$ is the position of jth atom in the $\gamma^{th}$ cell. k is a wave vector that may be chosen at the center or corner of the Brillouin zone. For $N$ cells in a system with $l$ atoms per cell, labeled by $x$ = 1, 2 . . . , $l$; the S operator from Equation 2 is written as:
@@
Here, $m,n$ are the indices of the phonon modes; $\hat{a}_{m}^{\dagger}$ and $\hat{a}_{n}$ are the creation and annihilation operators of a phonon mode respectively; ${\bf v}^{\bf k}_{mn}$ is the group velocity vector of mode $m$ when $m=n$; $\omega_{m}^{\bf k}$ is the vibrational frequency of the mode $m$. Upon applying the harmonic approximation results on Equation 1, a general expression for the real part of TC is obtained [12]:
@@
where, $f_{m}$ is the equilibrium occupation of the $m^{th}$ mode (i.e., the Bose-Einstein phonon distribution); $(S^{\alpha})_{mn}$ is the $\alpha^{th}$ component of $S_{mn}$. When $\omega\rightarrow 0$, due to the presence of the $\delta$ function, the term $\frac{\big{<}f_{m}\big{>}-\big{<}f_{n}\big{>}}{\hbar(\omega_{m}-\omega_{n})}$ in Equation 6 simplifies to $-{\partial\big{<}f_{m}\big{>}}/{\partial\omega_{m}}$. With these considerations, and using Equations 5, Allen and Feldman’s expression for the TC is:
@@
where, ${R}_{\gamma}^{\eta}$ is the $\eta$ component of position vector for cell $\gamma$; ${R}_{xx^{\prime}}^{\eta}$ is the $\eta$ component of displacement between atoms $x$ and $x^{\prime}$; $e^{\alpha,m}_{x,{\bf k}}$ is the polarization of the vibrational mode $m$ at atom $x$ along $\alpha$ direction. The thermal conductivity in Equation 2, is taken as an average of the Cartesian components of the conductivity tensor. It is worth noting that in the Allen-Feldman method, the lattice is properly quantized and so the method is suitable for arbitrarily low temperatures.
2.1 The Site-Projected Thermal Conductivity
@@
Rearranging summation order in Equation 2 and taking ${\bf k}={\bf 0}$, we can obtain an expression for $\kappa$ as a (double) sum over terms depending on atomic positions $x$ and $x^{\prime}$:
@@
For $\omega=0$, $\Xi$ is a Hermitian matrix, the thermal conduction matrix, i.e., $\Xi(x,x^{\prime})=\Xi(x^{\prime},x)$. In computing $\Xi$, care should be taken to exclude very small $\omega^{\bf{0}}_{i}$. Equation 8 suggests a way to decompose the total TC into contributions depending on atomic position $x$ by conducting a summation of $\Xi(x,x^{\prime})$ over all positions $x^{\prime}$ in the supercell:
@@
Here, $\zeta(x)$ can be named the site-projected thermal conductivity (SPTC), as it represents the contribution of an atom at site $x$ to the total TC of the system. Summing $\zeta(x)$ over atom index $x$ recovers the total TC $\kappa$:
@@ 3 Illustration of the Method
As a proof of concept, we conducted SPTC analysis on a range of different material systems, primarily carbon, silicon, and silicon-germanium structures. These systems included crystalline graphene, with both single and double vacancy defects, as well as amorphous graphene, characterized by ring disorder. Additionally, we applied SPTC analysis to diamond-crystalline silicon, para-crystalline silicon, and silicon-germanium alloy systems, focusing on grain boundaries and mass defects.
Molecular dynamics (MD) simulations were conducted using the Large-scale Atomic/Molecular Massively Parallel Simulator (LAMMPS) [14]. Within LAMMPS, the Machine Learning Gaussian approximation potential (ML-GAP) for carbon [15] and silicon [16] were employed for the carbon and pure silicon simulations. The Tersoff potential was used for the silicon-germanium structure [17].
@@ 3.2 Silicon Single-Vacancy Defect
A diamond-structure Si unit cell was sourced from the Materials Project [26], and a 1000-atom, $5\times 5\times 5$ supercell model was constructed from it and used as the base for creating different Si structures. The original cubic supercell, with a length of 27.21 Å  and density ($\rho$) of 2.33 g/cm-3, was relaxed, exhibiting no significant change in its conformation. A single-vacancy defect, shown in Figure 2(a), was created by removing an atom from the crystalline Si lattice, followed by another energy/structural relaxation. For the analysis of the temperature and frequency dependence of TC in amorphous silicon (aSi), a tetrahedral Si model with 512 atoms ($\rho=2.34$ g/cm-3), obtained from Reference [27], was utilized.
@@ 4.1 Carbon Structures
The SPTC for three-fold coordinated hexagonal rings in crystalline graphene are uniformly distributed. However, introducing a single-atom vacancy in an otherwise crystalline graphene structure significantly reduces the TC of nearby atoms. The values for the two-fold coordinated carbon atoms near the vacancy are approximately 27% that of the three-fold coordinated atoms in the bulk region.
As shown in Figure 3(a), the atoms adjacent to the vacancy (highlighted in blue) exhibit the lowest SPTC. An incremental, radially isotropic increase in normalized SPTC is observed moving outward from the defect center. To quantify anisotropy, the SPTC along two angular directions — 0 and 30, marked by the red and green lines in Figure 3(a) — is shown in Figure 3(b). The variation in SPTC with distance follows a similar trend in both directions.
@@
In the defective SW-585 configuration shown in Figure 3(c), the atoms in the five- and eight-membered rings are the lowest thermal conduction active sites, with the normalized SPTC reduced by approximately 60% compared to the bulk maximum. The SPTC gradually increases in all directions moving away from the defective region. The changes along two angular directions, 90 and 30, represented by red and green lines in Figure 3(c), are illustrated in Figure 3(d). The plot shows an isotropic increment in SPTC, returning to the maximum (bulk) values at around 10 Å  away from the defect center.
@@
In the aGr2000 model shown in Figure 4[i], atoms with only two neighbors exhibit SPTC values up to 74% of those of three-coordinated atoms. However, the SPTC for the three-coordinated carbon atoms varies depending on their local ring connectivity: atoms in non-hexagonal rings are less thermally active than those in hexagonal rings, similar to what is observed in crystalline graphene. To confirm this, a 60-atom amorphous graphene model (aGr60) was created. Its SPTC analysis, shown in Figure 4[ii], also indicates higher values for hexagonal rings compared to non-hexagonal ones. The ring distribution for both amorphous graphene models is illustrated in Figure 4[iii].
We note a similarity between thermally and electronically conductive regions. References [19] and [20], which report the electronic transport properties of amorphous graphite, report that in the amorphous graphene layers, the hexagonal carbon rings favor electron flow, while connections between pentagonal and heptagonal rings impede it. A similar observation was made for ring disorder in amorphous carbon nanotubes [29].
@@ 4.2 Silicon Structures
To test our implementation of Allen and Feldman formulation for the thermal conductivity [12], we compared the temperature and frequency dependence of TC in amorphous silicon produced by our implementation with the results of Feldman et al. [30], as shown in Figure 5(a) and (b). The TC temperature dependence obtained here is consistent with that in Reference [30], while the peak in the frequency dependence curve (the gray dashed line at 0.24 meV) matches the position they reported at 175 K. Note that the consistency is strong despite the use of a different interatomic potential and model in our work.
@@
In a defect-free diamond-silicon structure, all atoms contribute equally to the total TC of the system. On the other hand, local TC on atomic sites in the vicinity of the single-vacancy defect is reduced. Such under-coordinated silicon atoms with dangling bonds exhibit $\approx$70reduction in SPTC, as compared to the bulk-crystal value. The SPTC evolution with distance from the center of the defect in the (111) plane — along green and red lines shown in Figure 6(a) — is presented in Figure 6(b).
@@
In the SPTC analysis of the $\Sigma 3(112)/[1\bar{1}0]$ silicon grain boundary in Figure 7(a), the five-coordinated atoms (floating bonds) at the grain boundary exhibit lower SPTC compared to the bulk silicon and also influences the thermal activity of nearby atomic sites. Furthermore, a strong correlation between bonding environment and SPTC is evident: in Figure 7(b), the five-coordinated Si atoms labeled “i” show an SPTC reduction of around 79% compared to the bulk crystalline region labeled “ii.” Figure 7(c) further indicates that bond length distortions at the grain boundary contribute to the reduced SPTC, as atomic pairs with bond lengths exceeding 2.36 Å (the bond length in diamond silicon) exhibit significantly lower SPTC values.